SpaceRef

SpaceRef


Aqua Daily Report May 14, 2002 1130 GMT - May 15, 2002 1130 GMT

Status Report From: Goddard Space Flight Center
Posted: Wednesday, May 15, 2002

The following activities were completed:

  1. Uplinked the daily Master Command Loads and Ephemeris.
  2. The exercising of the Spacecraft 911 signal and path was successful, with the 1k telemetry data received at WSC and forwarded to the EOC for processing and display.
  3. The HSB instrument was successfully powered ON and placed in normal scan mode (1820z). HSB science data is now being collected and stored to the SSR.
  4. Loaded and dumped the SCSs in preparation for the GNC Sensor Calibrations. A negative 25 degree Yaw maneuver was performed first (maneuver start 0143z), followed by a positive 25 degree Yaw maneuver (maneuver start 0457z). The maneuvers are being performed as part of the IRU calibration.
  5. The CERES diagnostic memory patches to both the Fore and Aft instruments were loaded and dumped in preparation for the instrument brake release (0932z).
  6. Both the CERES Fore and Aft instrument brakes were successfully released (1045z). The Azimuth scan has transitioned from the ‘Caged’ position to the Cross Track position, a positive milestone in preparation for instrument on-orbit operations.
  7. Science data processing status: HSB Level 0 science data has been received by the JPL DAAC. Processing of the data to Level 1A will be performed today.

The following activities are scheduled for the next 24 hours:

  1. Load SCSs for the two remaining GNC Sensor Calibrations to be performed today. A positive 25 degree Pitch maneuver will be performed first (maneuver start 1445z), followed by a negative 25 degree Yaw maneuver (maneuver start 1945z). Roll maneuver “pairs” are scheduled during the evening shift (0219z, 0534z).
  2. The primary purpose of the GNC calibrations maneuvers are to calibrate the GNC sensors which need pointing accuracy for science operations.

Aqua Spacecraft Status:

  1. Spacecraft State is currently Standby.
  2. GNC Mode is Fine Point Mode, HiFi Ephemeris.

Aqua Instrument Status:

  1. AIRS is in Decontaminate state, Quiet bus is ON, Noisy bus is ON, Cryocooler is OFF, Scan mirror rotating is OFF and earth shield is CLOSED.
  2. AMSR-E is in Modified Sleep mode, rotating at 4 RPM.
  3. AMSU A1 is in Full Scan mode.
  4. AMSU A2 is in Full Scan mode.
  5. CERES Aft is in Safe mode and the covers are CLOSED.
  6. CERES Fore is in Safe mode and the covers are CLOSED.
  7. HSB is in Scan Normal mode and the antenna is in the TARGET position.
  8. MODIS is in Standby mode, Nadir Door is Unlatched and CLOSED, Solar Diffuser Door is Unlatched and CLOSED, and the Space View Door is Unlatched and CLOSED.

Activities Deferred:

  1. None.

Anomalies:

  1. Continuing to evaluate the apparent misalignment of approximately 0.2 degrees between Star Tracker Assembly (STA) 1 and STA2. A review of the mechanics associated with translating alignment measurement data into rotation matrices has been completed. No consequential discrepancies were found, working with Ball this week, to review the Star Tracker alignment data and conventions. The actions stated in the May 10 daily continue.
  2. There have been a total of 21 occurrences (3 new events uncovered yesterday from back-orbit data analysis) of the Command Telemetry Controller (CTC) receiving incomplete transfer frames from Transponder Interface Electronics (TIE) – (formerly CLCW errors not associated with commands). This anomaly has still not affected the ability to command the spacecraft and it has been safe at all times. The actions stated in the May 10 daily continue.
  3. A patch has been developed to address the timing issue between the CTC and the USO. Initial indications from testing with the ETSF were positive. The patch will be loaded as soon as final reviews are completed.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.